# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v2

@@ III Guided-DaSH
In this section, we describe our adaptation to the DaSH framework, Guided-DaSH, and explain the modifications we make to support motion planning problems without a highly structured search space via the use of guidance and to support problems with kinodynamic constraints. Figure 3 shows an overview comparison of Guided-DaSH and the original DaSH framework which we will elaborate on below. Algorithm 1 shows an overview of our framework.
@@
1:Guiding space $G$, Query $\{Q_{start},Q_{goal}\}$
@@
3:$C\leftarrow\emptyset$triangleright Guided search constraints
@@
5:    $\Pi^{G}\leftarrow\textsc{GuidedSearch}(G,C,Q_{start},Q_{goal})$
6:    $\mathcal{H}_{ts}\leftarrow\textsc{ConvertToTaskSpaceHypergraph}(\Pi^{G})$
7:    $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$triangleright $\mathcal{X}$ is an ordered list of vertices and hyperarcs in $\mathcal{H}_{ts}$
8:    $\mathcal{H}_{m}\leftarrow\emptyset$triangleright Initialize motion hypergraph
9:    for $x_{ts}\in\mathcal{X}_{ts}$
10:         $\pi\leftarrow\emptyset$triangleright Initialize local path
11:         if $X_{ts}.\textsc{IsVertex}$
12:             $\pi\leftarrow\textsc{ComputePath}(G,X_{ts})$
13:         else
14:             $\pi\leftarrow\textsc{ComputeTransition}(G,X_{ts})$
15:         if $\pi=\emptyset$
16:             $C\leftarrow C\cup\textsc{GetConstraint}(X_{ts})$
17:             break $\triangleright$ Restart guided search
18:         $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$triangleright Get elements in collision
19:         if $\mathcal{X}_{conf}\neq\emptyset$
20:             $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{X_{ts}\})$
21:             break $\triangleright$ Restart guided search
22:         else
23:             $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
24:    $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
III-A Guided High-level Search
In order to provide structure to a problem, the Guided-DaSH framework requires the identification of a guiding space $G$. To provide useful guidance, $G$ must reflect the structure of the planning space and should be quick to search. For example, in Section IV we propose a method that uses a topological skeleton of the workspace to provide guidance for mobile robot MRMP. Other possible guiding spaces include databases of multi-robot path segments or a discretization of each robot’s $C_{space}$. Identifying this space often requires domain knowledge of the problem being considered.
Each robot configuration should be able to be mapped to a state in $G$. Thus, when given a query $Q=\{Q_{start},Q_{goal}\}$ where $Q_{start}$ and $Q_{goal}$ are composite configurations containing start and goal values for each robot, the guided search returns a set of paths $\Pi^{G}=\{\pi^{G}_{r_{1}},...,\pi^{G}_{r_{n}}\}$ where each element of the paths is a state in $G$. The first element of each path is the state in $G$ corresponding to the robot’s start configuration and the last element corresponds to its goal. These paths then represent an abstract route that each robot will follow. The query may also be given a set of constraints specifying robots that cannot exist in the same state(s) together. These constraints are informed by the construction of the motion hypergraph and will be described in more detail in Section III-C.
States in $G$ should be defined at a resolution such that robots in the same state likely require coordination to avoid collision or accomplish their task. Using this knowledge, we can leverage $\Pi^{G}$ to construct the task space hypergraph.
III-B Task Space Hypergraph Construction
(a) Abstract paths over $G$
(b) Task Space Hypergraph
Figure 4: Abstract paths over the guiding space $G$ (a) and the constructed task space hypergraph (b) for a 4-robot group. Different colored blocks represent different states in $G$. As robots move into the same state in $G$, they are coupled into the same task space element in the task space hypergraph. Task space elements contain a set of robots as well as the constraints associated with the associated state in $G$.
The task space hypergraph $\mathcal{H}_{ts}=\{\mathcal{V}_{ts},\mathcal{E}_{ts}\}$ captures when coordination between robots is necessary. The original DaSH framework [5] exhaustively constructs the task space hypergraph; however, if we use guidance to decide in advance which task space elements are necessary, we can efficiently construct a minimal portion of the task space hypergraph. This makes our adapted framework advantageous for problems where exhaustive construction is computationally intractable. Given the abstract paths for each robot $\Pi^{G}$, we can translate the sequence of states in $G$ that each robot visits into task space elements $\mathcal{V}_{ts}$ and transitions between them $\mathcal{E}_{ts}$.
Algorithm 2 shows an overview of the task space hypergraph construction procedure. We begin by creating a virtual source $V_{source}$ that denotes the start of the problem. Then, at each timestep in the abstract paths, we group together robots in the same state in $G$ and add vertices to the task space hypergraph reflecting that robot group and the constraints associated with the corresponding state in $G$. Transitions between the previous timesteps and current timesteps are also added. These hyperarcs capture both changes in composition and constraints of the task space elements they connect. Finally, the virtual target is added denoting the end of the problem and connected via a hyperarc to the last task space element for each robot. Figure 4 shows an example of this process.
Algorithm 2 Task Space Hypergraph Construction
1:Abstract paths $\Pi^{G}=\{\pi^{G}_{r_{1}},...,\pi^{G}_{r_{2}}\}$
2:$T\leftarrow\max_{r\in R}\|\pi^{G}_{r}\|$triangleright Maximum path timesteps
3:$\mathcal{H}_{ts}\leftarrow\{V_{source},\emptyset\}$triangleright Initialize task space hypergraph
4:$S_{prev}\leftarrow\{(V_{source},R)\}$triangleright Previous $G$ state of each robot
5:for $t\in\{1,...,T\}$
6:    $S_{cur}\leftarrow\emptyset$triangleright Current $G$ state of each robot
7:    for $r\in R$
8:         $S_{cur}[\pi^{G}_{r,t}]\leftarrow S_{cur}[\pi^{G}_{r,t}]\cup\{r\}$
9:    for $\{state,robots\}\in S_{cur}$
10:         $\mathcal{H}_{ts}.\textsc{AddVertex}(\{state,robots\})$
11:    $\mathcal{E}_{ts}\leftarrow\textsc{GetTransitions}(S_{prev},S_{cur})$
12:    $\mathcal{H}_{ts}.\textsc{AddHyperarcs}(\mathcal{E}_{ts})$
13:    $S_{prev}\leftarrow S_{cur}$
14:$\mathcal{H}_{ts}.\textsc{AddVertex}(V_{target})$triangleright Connect last vertices to target
15:$\mathcal{H}_{ts}.\textsc{AddHyperarc}(S_{prev},V_{target})$
III-C Motion Hypergraph Construction
@@
The motion hypergraph $\mathcal{H}_{m}$ stores the configurations and motions between them that make up the robots’ motion paths as well as the composition and transition hyperarcs from $\mathcal{H}_{ts}$. The abstract paths in $G$ from which $\mathcal{H}_{ts}$ is constructed are only meaningful if the robots’ motion paths actually follow the sequence of states $\Pi^{G}$. Thus, it is often necessary to leverage a guided planning method to ensure that the constraints associated with each task space element are met when constructing local paths. We detail an example of such as method in Section IV.
@@
For each task space element, a local path is computed in the composite space of the contained therein. These paths become vertices and motion hyperarcs in $\mathcal{H}_{m}$. Transition hyperarcs in $\mathcal{H}_{ts}$ must also be translated into local paths. Transitions often require a different method of path construction to account for, e.g., grasp constraints when picking up an object or other factors when moving between states in $G$.
It is possible that constructing a local path for a task space element or transition could fail. In this case of a task space element failing, a constraint is added to the guided search to prohibit that group of robots from existing in the corresponding state in $G$ together. Similarly, in the case of a transition failing, constraints are added to prevent the groups of robots in the tail and head sets of the transition from being in the corresponding incoming and outgoing states in $G$ together.
@@
In the original DaSH formulation, the entire motion hypergraph is constructed before a scheduled query is performed that finds and resolves conflicts between robot paths through different task space elements or transitions. In our adapted framework; however, since we know from the guided search which task space element’s motion should occur concurrently, we can interleave the motion hypergraph construction and conflict detection stages. This allows us to reduce wasted computation by finding conflicts as they occur rather than after the motion hypergraph is constructed which would likely result in local paths being constructed that would not be included in the final paths.
III-C1 Kinodynamic Planning
@@
III-C2 Conflict Detection
As we construct each movement hyperarc corresponding to motion through a task space element, we track the high-level timestep that this segment of the motion hypergraph corresponds to in the guided search. Additionally, when constructing the movement hyperarcs corresponding to transitions between task space elements, we classify these local paths as extensions of the local paths corresponding to the tail set of task space elements. Then, after each local path is computed, it is checked for collision against each existing local path occurring at the same high-level timestep by considering fine-resolution intermediate configurations along the paths.
When a conflict is found, constraints are added to the guided search to prevent the robots in each task space element or transition from existing in or between the corresponding states in $G$ together. When new constraints are added, the guided search is redone, the task space hypergraph is recomputed, and the construction of the motion hypergraph restarts.
III-D Query
The original DaSH framework involves a query stage after the construction of the motion hypergraph to find an optimistic and then a validated schedule of motions for the robots to follow. This is necessary under exhaustive construction of the task space and motion hypergraphs; however, because Guided-DaSH only constructs the portion of the task space hypergraph needed to find a motion solution guided by the abstract paths and resolves conflicts during construction of the motion hypergraph, this stage is no longer necessary. Instead, we can directly extract a motion path for the robots from the computed motion hypergraph and terminate.
@@ IV Workspace Guided-DaSH
In this section, we explain how we apply the Guided-DaSH framework for mobile robot MRMP in congested environments with narrow passages to propose the Workspace Guided-DaSH method and describe how we leverage the workspace skeleton representation of the environment to quickly find motion plans for large robot teams. The workspace skeleton functions as the guiding space $G$ and directs both the construction of the task space hypergraph and its translation into the motion hypergraph.
Algorithm 3 Workspace Guided-DaSH
1:Skeleton $S$ and a Query $\{Q_{start},Q_{goal}\}$
2:$\Pi\leftarrow\emptyset$triangleright Robot motion paths
3:$C\leftarrow\emptyset$triangleright MAPF search constraints
4:while $\Pi=\emptyset$
5:    $\Pi^{G}\leftarrow\textsc{MAPF}(S,C,Q_{start},Q_{goal})$
6:    $\mathcal{H}_{ts}\leftarrow\textsc{ConvertToTaskSpaceHypergraph}(\Pi^{G})$
7:    $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$triangleright $\mathcal{X}$ is an ordered list of vertices and hyperarcs in $\mathcal{H}_{ts}$
8:    $\mathcal{H}_{m}\leftarrow\emptyset$triangleright Initialize motion hypergraph
9:    for $x_{ts}\in\mathcal{X}_{ts}$
10:         $\pi\leftarrow\emptyset$triangleright Initialize local path
11:         if $X_{ts}.\textsc{IsVertex}$
12:             $\pi\leftarrow\textsc{SimplifiedCDRRRT}(S,X_{ts})$
13:         else
14:             $\pi\leftarrow\textsc{CompositeRRT}(X_{ts})$
15:         if $\pi=\emptyset$
16:             $C\leftarrow C\cup\textsc{GetConstraint}(X_{ts})$
17:             break $\triangleright$ Restart guided search
18:         $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$triangleright Get elements in collision
19:         if $\mathcal{X}_{conf}\neq\emptyset$
20:             $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{X_{ts}\})$
21:             break $\triangleright$ Restart guided search
22:         else
23:             $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
24:    $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
Algorithm 4 SimplifiedCDRRRT
1:Task space element $V$, maximum extension failures $\tau$
2:$T\leftarrow\emptyset$triangleright Initialize empty tree
[…]
4:$coveredEdge\leftarrow false$
5:while $\neg coveredEdge$ and $r.\textsc{NumFailures}<\tau$
6:    $q_{rand}\leftarrow\textsc{SampleCfg}(r)$triangleright Sample from region
7:    $q_{new}\leftarrow\textsc{ExtendLocalPath}(T,q_{rand})$
[…]
10:    else
11:         while $r.\textsc{ContainsCfg}(q_{new})$
12:             $r.\textsc{AdvanceAlongSkeletonEdge}(V.\textsc{edge})$
13:             if $r.\textsc{ReachedEndOfSkeletonEdge}(V.\textsc{edge})$
14:                 $coveredEdge\leftarrow true$
15:if $r.\textsc{NumFailures}>\tau$
16:    return $\emptyset$
17:return $T$
@@ IV-A Task Space Hypergraph Construction
The task space hypergraph captures when sets of robots should be composed into the same planning space. To construct a meaningful task space hypergraph, we must define the task space elements in a way that captures when coordination between mobile robots is needed. Considering mobile robots, coordination between the motions of robots is required when they are physically in proximity to each other. Our insight is that we can leverage the workspace skeleton to approximate when coordination is necessary by modeling the movement of the robots through the workspace as movement along the skeleton. Specifically, coordination is necessary when robots are traversing the same skeleton edge (either in the same or opposite direction) or traveling through the same skeleton vertex. A task space element then contains a set of robots and the corresponding segment of the skeleton around which planning is constrained. Transitions model sets of robots moving through the same skeleton vertex. This provides structure to the problem, allowing us to decompose the task space.
Exhaustively constructing the task space hypergraph, which has previously been done for DaSH methods [5], is intractable for the case we consider because the structure we induce on the task space does not allow for as sparse of a task space hypergraph representation as seen in the previously considered problems. Using the Guided-DaSH framework, we instead construct this representation minimally and as needed by exploiting the guidance provided by the skeleton. We use MAPF to generate abstract paths for the robots to move over the skeleton from the vertex closest to their start to the vertex closest to their goal.
Algorithm 3 shows an overview of Workspace Guided-DaSH. Given a skeleton of the workspace, we first construct a path over the skeleton for each robot using a MAPF algorithm (line 5). We set the capacity of each skeleton vertex and edge as given by the minimum clearance to an obstacle. We then combine these paths to form the minimal necessary portion of the task space hypergraph (line 6). Fig. 5 shows an example of the conversion of the MAPF solution into the task space hypergraph. Robots traversing the same skeleton edge are grouped into the same task space element using composition hyperarcs. As they move away from each other onto different edges, they are decoupled into separate task space elements, boosting planning efficiency by considering smaller $\mathcal{C}_{spaces}$. Transition hyperarcs capture robots moving through the same skeleton vertex, providing the necessary coordination, as well as robots moving between segments of the skeleton (i.e. moving between task space elements with the same robots but different constraints). We then use the task space hypergraph to guide the construction of the motion hypergraph.
@@
Figure 5: A segment of the robot paths over the skeleton and the corresponding portion of the task space hypergraph. Each task space element shows the corresponding set of robots and the relevant portion of the topological skeleton. In (b), arrows indicate robots moving along a skeleton edge and dots indicate a decoupled waiting state. Solid lines indicate transition hyperarcs and dashed lines represent composition hyperarcs. The green and red robots are initially moving in opposite directions along the same skeleton edge. Then, red and blue reach the skeleton vertex together and go on to different edges. The yellow robot begins moving in the opposite direction of the red robot when the red robot reaches the next skeleton edge.
@@ IV-B Translation to Low-level Motion Hypergraph
Once the task space hypergraph is constructed, we compute a set of paths to form the motion hypergraph. The task space hypergraph is only meaningful if the robots roughly follow their abstract paths over the skeleton. We must also traverse narrow passages in the environment. Thus, we leverage the skeleton to guide the construction of local paths along skeleton edges. To compute paths through task space elements, we use a simplified variant of CDR-RRT (Sec. II-D2; Alg. 3 line 12; Alg. 4) that advances a dynamic region along the skeleton edge.
Considering transition hyperarcs that represent multiple robots moving through a skeleton vertex, we stop the movement of the robots along the incoming skeleton edges a small distance $\delta$ from the end of each edge. We’ve found that setting $\delta$ to the diameter of the region used for CDR-RRT [10] works well. We then grow an RRT [14] in the composite space of these robots from the last configuration on the robots’ incoming skeleton edges to a sampled configuration at least $\delta$ distance away from the skeleton vertex on the robots’ next skeleton edges (Alg. 3, line 14). Here, we simply grow an RRT rather than using CDR-RRT to avoid forcing the robots to reach the skeleton vertex at the same time, creating congestion. Waiting vertices, composition hyperarcs, and transition hyperarcs not associated with movement in the task space hypergraph correspond directly to vertices and hyperarcs in the motion hypergraph.
It is possible for construction of a local path to fail if the RRT fails to be extended a given number of times during planning. In this case, we impose a constraint that the MAPF solution from which we build the task space hypergraph cannot contain the same group of robots traversing along either a skeleton edge that failed or the same groups of robots traversing along the incoming and outgoing skeleton edges to a skeleton vertex that failed (line 16). We then recompute the task space hypergraph and try again to construct a complete motion hypergraph. Similarly, if conflicts are found between robots traversing different portions of the skeleton simultaneously, constraints are imposed to prevent the relevant robot groups from traversing these segments of the skeleton together (line 20).

### v3

@@ III Guided-DaSH
In this section, we describe our adaptation to the DaSH framework, Guided-DaSH, and explain the modifications we make to support motion planning problems without a highly structured search space and kinodynamic constraints. For a detailed description of the DaSH framework, see [5]. Figure 1 shows a comparison of Guided-DaSH and the original DaSH framework which we will elaborate on below. Algorithm 1 shows an overview of our framework.
@@
1:Heuristic $H$, Query $Q$
@@
3:$C\leftarrow\emptyset$triangleright Heuristic search constraints
@@
5:    $\mathcal{H}_{ts}\leftarrow\textsc{ConstructTaskSpaceHypergraph}(H,C,Q)$
6:    $\mathcal{X}_{ts}\leftarrow\textsc{OrderByDependency}(\mathcal{H}_{ts})$
7:    $\mathcal{H}_{m}\leftarrow\emptyset$triangleright Initialize motion hypergraph
8:    for $x_{ts}\in\mathcal{X}_{ts}$
9:         $\pi\leftarrow\textsc{ComputePath}(G,x_{ts})$
10:         if $\pi=\emptyset$
11:             $C\leftarrow C\cup\textsc{GetConstraint}(x_{ts})$
12:             break $\triangleright$ Restart heuristic search
13:         $\mathcal{X}_{conf}\leftarrow\textsc{FindCollisions}(\mathcal{H}_{m},\pi)$
14:         if $\mathcal{X}_{conf}\neq\emptyset$
15:             $C\leftarrow C\cup\textsc{GetConstraints}(\mathcal{X}_{conf}\cup\{x_{ts}\})$
16:             break $\triangleright$ Restart heuristic search
17:         $\mathcal{H}_{m}.\textsc{AddPath}(\pi)$
18:    $\Pi\leftarrow\textsc{ExtractMotionPaths}(\mathcal{H}_{m})$
III-A Task Space Hypergraph Construction
In the task space hypergraph $\mathcal{H}_{ts}=\{\mathcal{V}_{ts},\mathcal{E}_{ts}\}$, vertices $\mathcal{V}_{ts}$ are task space elements which represent local planning spaces and are composed of the robots required to coordinate at that point and relevant constraints they should adhere to. Hyperarcs $\mathcal{E}_{ts}$ can be categorized as composition, representing changes in the robots considered and constraints, or transition, representing changes in constraints for a fixed set of robots. The original DaSH framework [5] exhaustively constructs $\mathcal{H}_{ts}$; however, by using guidance to inform which task space elements are needed, we can efficiently construct a minimal portion of $\mathcal{H}_{ts}$. This makes our adapted framework advantageous for problems where exhaustive construction is intractable.
To construct a minimal portion of the task space hypergraph, we require guidance in the form of domain knowledge about the problem being considered. This defines the set of constraints which can exist in a task space element. Additionally, this knowledge must inform a heuristic $H$ required to search over the implicit problem space and construct a series of task space elements $\mathcal{V}_{ts}$ that satisfy a given query. Hyperarcs $\mathcal{E}_{ts}$ are added between subsequent task space elements to form $\mathcal{H}_{ts}$. The heuristic must also take as input a set of constraints that represent prohibited task space elements that cannot exist in the solution since they have failed to be translated into a motion solution. Searching over the implicit problem space can be achieved by searching over a smaller representative space that captures when coordination between robots is required. We provide an example in which information about the workspace topology is used to guide composition decisions in Section IV.
III-B Motion Hypergraph Construction
@@
The motion hypergraph $\mathcal{H}_{m}$ stores the configurations and motions between them that make up the robots’ motion paths through the task space elements and hyperarcs from $\mathcal{H}_{ts}$. It is often necessary to leverage a guided planning method to ensure that the constraints associated with each task space element are met when constructing local paths. We detail an example of such a method in Section IV.
@@
For each task space element and transition hyperarc in $x_{ts}\in\mathcal{H}_{ts}$, ordered from start to goal, a local path is computed in the composite space of the robots contained therein. These paths become vertices and motion hyperarcs in $\mathcal{H}_{m}$. Transitions often require a different method of path construction to account for, e.g., grasp constraints or other problem-specific factors. It is possible that constructing a local path for a task space element or transition could fail. In this case of a task space element failing, a constraint is added to the heuristic search to prohibit that task space element from existing in $\mathcal{H}_{ts}$. Similarly, in the case of a transition failing, constraints are added to prohibit the task space elements in the corresponding incoming and outgoing sets of the transition from existing simultaneously in $\mathcal{H}_{ts}$.
@@
In the original DaSH formulation, the entire motion hypergraph is constructed before a scheduled query is performed that finds and resolves conflicts between robot paths through different task space elements or transitions. In our adapted framework; however, since we know from the heuristic search which task space element’s motions should occur concurrently, we can interleave motion hypergraph construction and conflict detection. By finding conflicts as they occur rather than after the motion hypergraph is constructed, we minimize wasted computation and can directly extract the motion solution from the motion hypergraph, bypassing the query stage.
@@
As we construct $\mathcal{H}_{m}$, after each local path is computed, we check for collisions with existing local paths occurring simultaneously by considering intermediate configurations along the paths. When a conflict is found, constraints are added to the heuristic search to prevent the corresponding task space elements or transitions from existing in $\mathcal{H}_{ts}$ together. When constraints are added, the heuristic search is redone, $\mathcal{H}_{ts}$ is recomputed, and construction of $\mathcal{H}_{m}$ restarts. This process continues until a conflict-free motion solution is found.
@@ IV Workspace Guided-DaSH
In this section, we explain how we apply the Guided-DaSH framework for mobile robot MRMP in congested environments with narrow passages to propose the Workspace Guided-DaSH method and describe how we leverage the workspace skeleton to guide both the construction of $\mathcal{H}_{ts}$ and its translation into $\mathcal{H}_{m}$.
@@ IV-A Task Space Hypergraph Construction
Considering mobile robots, coordination between the motions of robots is required when they are physically in proximity to each other. Our insight is that we can leverage the skeleton to approximate when coordination is necessary by modeling the movement of the robots through the workspace as movement along the skeleton. Specifically, coordination is necessary when robots are traversing the same skeleton edge (either in the same or opposite direction) or traveling through the same skeleton vertex. This formulation imposes a structure on the search space that we capture through our heuristic search.
Workspace Guided-DaSH follows the structure shown in Alg. 1. Given a skeleton of the workspace, we first conduct our heuristic search to find a path over the skeleton for each robot using a MAPF algorithm. We set the capacity of each skeleton vertex and edge as given by the minimum clearance to an obstacle. We then combine these paths to form the minimal necessary portion of the task space hypergraph. Fig. 2 shows an example of the conversion of the MAPF solution into $\mathcal{H}_{ts}$. Within task space elements, constraints represent the skeleton segments that each robot should traverse at that stage. Robots traversing the same skeleton edge are grouped into the same task space element using composition hyperarcs. As robots move away from each other onto different edges, they are decoupled into separate task space elements, boosting planning efficiency by considering smaller $\mathcal{C}_{spaces}$. Transition hyperarcs capture robots moving through the same skeleton vertex as well as robots moving between segments of the skeleton (i.e. moving between task space elements with the same robots but different constraints).
@@
Figure 2: A segment of the robot paths over the skeleton (gray) and the corresponding portion of the task space hypergraph. Each task space element shows the corresponding set of robots and the relevant portion of the topological skeleton. In (b), arrows indicate robots moving along a skeleton edge and dots indicate a decoupled state. Solid lines indicate transition hyperarcs and dashed lines represent composition hyperarcs. The green and red robots are initially moving in opposite directions along the same skeleton edge. Then, red and blue reach the skeleton vertex together and go on to different edges. The yellow robot begins moving in the opposite direction of the red robot when the red robot reaches the next skeleton edge.
@@ IV-B Translation to Low-level Motion Hypergraph
Once $\mathcal{H}_{ts}$ is constructed, we compute a set of paths to form $\mathcal{H}_{m}$. To ensure that the robots roughly follow their paths over the skeleton and that narrow passages in the environment are efficiently traversed, we leverage the skeleton to guide the construction of local paths along skeleton edges. To compute paths through task space elements, we use the dynamic sampling region advancement procedure from CDR-RRT [6].
Considering transition hyperarcs that represent multiple robots moving through a skeleton vertex, we stop the movement of the robots along the incoming skeleton edges a small distance $\delta$ from the end of each edge. We’ve found that setting $\delta$ to the diameter of the region used for the dynamic sampling region advancement works well. We then grow an RRT [14] in the composite space of these robots from the last configuration on the robots’ incoming skeleton edges to a sampled configuration at least $\delta$ distance away from the skeleton vertex of the robots’ next skeleton edges. Here, we simply grow an RRT rather than using CDR-RRT to avoid forcing the robots to reach the skeleton vertex at the same time, creating congestion. Composition hyperarcs and transition hyperarcs not associated with movement in the task space hypergraph correspond directly to vertices and hyperarcs in the motion hypergraph.
We consider construction of a local path to have failed if the RRT fails to be extended a given number of times during planning. In this case, we impose a constraint that the MAPF solution from which we build $\mathcal{H}_{ts}$ cannot contain the same group of robots traversing along either a skeleton edge that failed or the same groups of robots traversing along the incoming and outgoing skeleton edges to a skeleton vertex that failed. We then recompute $\mathcal{H}_{ts}$ and try again to construct $\mathcal{H}_{m}$. Similarly, if conflicts are found between robots traversing different portions of the skeleton simultaneously, constraints are imposed to prevent the relevant robot groups from traversing these segments of the skeleton together.